# Quantum Metrology via Floquet-Engineered Two-axis Twisting and Turn Dynamics

### v1

@@ Acknowledgements.
We can further estimate the global optimal parameter $\delta_{\rm opt}$ and the explicit expression for the time evolution to reach maximum QFI using the semi-classical description. We find that the TAT-and-turn Hamiltonian can generate GHZ-like states, whose variance is stretched along $\alpha$-axis or $\beta$-axis on the $x-y$ plane. For calculation clarity, we start from the initial position $(0,0,1)$, when $\delta_{\mathrm{eff}}>0$ (default $\chi>0$), we take $\hat{J}_{\beta}$ as the generator and obtain
$F_{Q}=4\left(\Delta\hat{J}_{\beta}\right)^{2}=N^{2}B^{2}=\frac{N^{2}}{2}(1-Z)(1+Z+\frac{2\delta_{\mathrm{eff}}}{N\chi_{\mathrm{eff}}}).$ (39)
When $Z=\frac{\delta_{\mathrm{eff}}}{N\chi_{\mathrm{eff}}}$, the above Eq. (39) reaches its maximum about variable $Z$, and is converted into
$F_{Q}=\frac{N^{2}}{2}(1-\frac{\delta_{\mathrm{eff}}}{N\chi_{\mathrm{eff}}})^{2}\leq(F_{Q})_{\mathrm{max}}\leq N^{2}.$ (40)
One can obtain the maximum QFI $(F_{Q})_{\mathrm{max}}$ when
@@
And for $\delta_{\mathrm{eff}}<0$, the expression of QFI is $F_{Q}=4\left(\Delta\hat{J}_{\alpha}\right)^{2}=N^{2}A^{2}\leq N^{2}$ with the generator being $\hat{J}_{\alpha}$, the corresponding optimal detuning is ${\delta_{\rm opt}}/({N\chi_{\mathrm{eff}}})=1-\sqrt{2}\approx-0.414$.
We consider $\delta_{\mathrm{eff}}>0$ (the calculation process for $\delta_{\mathrm{eff}}<0$ is the same as below), within the time range $0\leq t\leq t_{\rm opt}$ of reaching the maximum QFI. We find that $AB=\frac{1}{2}(1-Z)\sqrt{(1+Z)^{2}-(\frac{2\delta_{\mathrm{eff}}}{N\chi_{\mathrm{eff}}})^{2}}>0$, and obtain a differential equation that depends only on variable $Z$, which is given by
@@
This result is obtained via the property of uncertainty patch. The lower bound $\sqrt{1-\frac{1}{2N}}$ of the integral comes from the separatrix of the uncertainty patch, which determines the initial stretching position, and the integral upper bound $\frac{2\delta_{\mathrm{eff}}}{N\chi_{\mathrm{eff}}}-1$ is the critical point of variance change on the separatrix. The Taylor expansion used above is
@@
$f[Z(t)]-f[Z(0)]+\sqrt{{(N\chi_{\mathrm{eff}}})^{2}-\delta_{\mathrm{eff}}^{2}}t=0,$ (47)
@@
Due to the special symmetry of the phase space of the TAT-and-turn Hamiltonian, the same result can be obtained for the another saddle point $(0,0,-1)$, the calculation process is similar to the above.
IV IV. Heisenberg scaling and comparison with spin cat states
Figure 5: (a) The probability distributions of non-Gaussian state generated by Floquet-engineered TAT-and-turn dynamics and its corresponding spin cat state with $\theta=\arccos{\overline{Z}}$ for particle number $N=100$. (b) The QFI of non-Gaussian states generated by Floquet-engineered TAT-and-turn dynamics and their corresponding spin cat states with $\theta=\arccos{\overline{Z}}$ versus different particle number $N$.
@@
In the main text, we have shown that the non-Gaussian states generated by Floquet-engineered TAT-and-turn dynamics have metrological properties of Heisenberg scaling, which are similar to the spin cat states. In this section, we further analyze the QFI scaling of the non-Gaussian states with optimal QFI generated using Floquet-engineered TAT-and-turn dynamic by approximating the generated non-Gaussian entangled states as spin cat states. Compared with using the semi-classical method of taking the critical point $Z(t_{\rm opt})\approx\frac{2\delta_{\mathrm{eff}}}{N\chi_{\mathrm{eff}}}-1$ through Eq. (39) in the previous section, the obtained results are closer to the numerical results using full quantum treatment.
@@
We consider the non-Gaussian state $\lvert\psi\rangle_{\rm NGS}$ generated by Floquet-engineered TAT-and-turn dynamic starting from the initial state $\lvert\uparrow\rangle^{\otimes N}$ with $\alpha=0$, the time in Eq. (46) and the critical detuning in Eq. (41), whose two main parts are symmetrically distribution near the pole of $y$-axis. To unify the orientation with a spin-cat state consisting of two SCSs with the same azimuthal angle $\varphi$ and polar angle symmetric about $\vartheta=\pi/2$, the non-Gaussian state $\lvert\psi\rangle_{\rm NGS}$ need to be rotated to the $z$-axis via $e^{-i\hat{J}_{x}\pi/2}$. The expectation of $\hat{J}_{z}$ for the non-Gaussian state $\langle\hat{J}_{z}\rangle_{\rm NGS}^{{}^{\prime}}=0$ with $\lvert\psi\rangle_{\rm NGS}^{{}^{\prime}}=e^{-i\hat{J}_{x}\pi/2}\lvert\psi\rangle_{\rm NGS}$, and the probability distribution is mainly concentrated around the two ends of the $z$-axis. Thus the symmetrical non-Gaussian entangle state can be approximately regarded as a macroscopic superposition spin coherent state (MSSCS) like a spin cat state [48], which is given by
@@
$\lvert\psi\rangle_{\rm NGS}^{{}^{\prime}}\sim|\Psi(\vartheta)\rangle_{\mathrm{CAT}}\approx\frac{1}{\sqrt{2}}(|\vartheta,\varphi\rangle_{\rm SCS}+|\pi-\vartheta,\varphi\rangle_{\rm SCS}),$ (48)
@@
where $\vartheta=\arccos\overline{Z}$ and the normalized average spin length $\overline{Z}=\frac{2}{N}\sum_{m=-N/2}^{N/2}|m|P_{m}=\frac{2}{N}\sum_{m=-N/2}^{N/2}|m||\langle m|\psi\rangle^{{}^{\prime}}_{\rm{NGS}}|^{2}$ with the eigenstate $|m\rangle$ of $\hat{J}_{z}$. It approximates the non-Gaussian state as a spin cat state with the same normalized average spin length $\overline{Z}$. The probability distribution of this non-Gaussian state and its corresponding spin cat state are shown in Fig. 5 (a). In this case, the QFI of this non-Gaussian state is approximately written as
$F_{Q}^{\rm NGS}\sim F_{Q}^{\rm CAT}\approx N^{2}\cos^{2}\vartheta=N^{2}{\overline{Z}}^{2},$ (49)
@@
Further, we use Eq. (49) to calculate the QFI of non-Gaussian states generated by TAT-and-turn dynamics for different particle number $N$ and find that the results are in agreement with the quantum results, as shown in Fig. 5 (b), illustrating that the best achievable QFI of Floquet-engineered TAT-and-turn exhibits the Heisenberg-limited scaling $F_{Q}^{\rm max}\propto N^{2}$, which is consistent with the QCRB of a spin cat state.
V V. Optimal first-order measurement
@@
For a final state after given dynamics with interaction-based readout or half-population difference readout, the achievable measurement precision depends both on the directions to encode signal into probe state and to detect the final state. The signal encoding operator $\hat{J}_{\vec{n}}$ determines the QFI of a given probe state, thus setting a bound (i.e., QCRB) on the ultimate measurement precision. The detection using observable operator $\hat{J}_{\vec{m}}$ determines how closely we can approach the QCRB from a given measurement direction. Optimization in both directions $\vec{n}$ and $\vec{m}$ is of great significance to improving measurement precision. The method to optimize the measurement direction with a given sensing direction is shown as follows.
Figure 6: The interaction-based readout (IBR) at $\phi=1/1000$ for (a) OAT, (b) OAT-and-turn, (c) TAT, and (d) TAT-and-turn with particles number $N=100$ at evolution time $t=t_{1}=t_{2}$. (e) The attainable measurement precision of TAT-and-turn at $t=0.12$ via IBR for $\phi$ approaching 0.
@@
In the general process of interaction-based readout, for an arbitrary observable operator $\hat{J}_{\vec{m}}$, we transform the Schrödinger picture into the Heisenberg picture, i.e.
@@
where $\hat{R}_{\vec{n}}(\phi)=e^{-i\hat{J}_{\vec{n}}\phi}$ for encoding signal, $\hat{U}_{1}=e^{-i\hat{H}_{1}t_{1}}$ for generating entangled states, and $\hat{U}_{2}=e^{-i\hat{H}_{2}t_{2}}$ for signal amplification through disentanglement, typically time-reversal IBR with $\hat{H}_{2}=-\hat{H}_{1}$ and $t_{2}=t_{1}$. When half-population difference readout is performed directly instead of interaction-based readout, the Hamiltonian $\hat{H}_{2}=0$. Therefore we can calculate measurement response and express it in matrix form similar to Eq. (24) or Ref. [95], which is given by
@@
where $\mathcal{Q}$ represents the $3\times 3$ covariance matrix with elements $\mathcal{Q}_{\mu\nu}=\langle\{\hat{\tilde{J}}_{\mu},\hat{\tilde{J}}_{\nu}\}\rangle/2-\langle\hat{\tilde{J}}_{\mu}\rangle\langle\hat{\tilde{J}}_{\nu}\rangle$. Based on the error propagation formula, the achievable measurement sensitivity becomes
@@
To find the optimal measurement direction, one can using the Cauchy-Schwartz inequality $(\vec{u}^{T}\vec{v})^{2}\leq\vec{u}^{T}\vec{u}\vec{v}^{T}\vec{v}$, the measurement sensitivity satisfies
@@
in which $\mathcal{K}=\mathcal{M}\mathcal{Q}^{-1}\mathcal{M}^{T}$ depends on the selection of $\vec{n}$, and the equality is achieved at $\vec{m}_{\rm opt}=\mathcal{N}_{c}\mathcal{Q}^{-1}\mathcal{M}^{T}\vec{n}$ with $\mathcal{N}_{c}$ the normalization constant, a given sensing operator $\hat{J}_{\vec{n}}$ for the general case including $\phi=0$ and $\phi\neq 0$.
Thus, after applying the optimal detection direction, i.e. $\vec{m}_{\rm opt}=\mathcal{N}_{c}\mathcal{Q}^{-1}\mathcal{M}^{T}\vec{n}$, the measurement precision only determined by the encoding signal direction $\vec{n}$. In this work, we choose the sensing direction $\vec{n}_{\rm max}$ corresponding to the maximum QFI value of the entangled state to encode the signal. Thus combining Eq. (27), Eq. (53) and Eq. (54), we arrive
@@
Note that the phase $\phi$ cannot be equal to 00, otherwise the matrix $\mathcal{K}$ will lose the direction $\vec{n}_{\rm max}$ and be restricted to the plane perpendicular to the polarization direction.
As shown in Fig. 6 (a-d), we analyze the measurement precision of several common nonlinear Hamiltonians such as OAT, OAT-and-turn, TAT, and our TAT-and-turn under time-reversal IBR using above method. The results show that the method we proposed is close to the QCRB with excellent performance of measurement precision. The entangled state generated in a short time is a spin squeezed state, and after a longer time, the generated state becomes a non-Gaussian state. During the process, the measurement precision can approach the QCRB. Especially for OAT, this method can saturate QCRB in most of the time range, which achieves ideal measurement precision. From the Fig. 6 (e), we find that the measurement precision gradually converges as $\phi$ approaches 00, which is conducive to high-precision measurement of tiny signals.
VI VI. Experimental feasibility of Floquet-engineered TAT-and-turn dynamics
In an interacting atomic ensemble driven by an external coupling field, the system obey the Hamiltonian $\hat{H}=\chi\hat{J}_{z}^{2}+\omega_{0}\hat{J}_{z}+\Omega(t)[e^{i(\omega_{c}t+\alpha)}+e^{-i(\omega_{c}t+\alpha)}]\hat{J}_{x}$, where $\omega_{0}$ is the atomic transition frequency, $\omega_{c}$ is the frequency of the coupling field, $\alpha$ is the phase of coupling field, and $\Omega(t)$ is the Rabi frequency. To realize the Hamiltonian (9), we choose to modulate the Rabi frequency $\Omega(t)=2\Omega_{0}\cos(\omega t)$, i.e.
$\hat{H}=\chi\hat{J}_{z}^{2}+\omega_{0}\hat{J}_{z}+2\Omega_{0}\cos\left(\omega t\right)\cos\left(\omega_{c}t+\alpha\right)\hat{J}_{x}$ (56)
In the rotating-frame with rotating-wave approximation (RWA), let $\hat{U}=e^{-i\omega_{c}t\hat{J}_{z}}$, the Hamiltonian becomes
$\hat{H}_{R}$ $=\hat{U}^{\dagger}\hat{H}\hat{U}-i\hat{U}^{\dagger}\frac{\partial\hat{U}}{\partial t}$ (57)
$=\chi\hat{J}_{z}^{2}+\omega_{0}\hat{J}_{z}+e^{i\omega_{c}t\hat{J}_{z}}\left[2\Omega_{0}\cos\left(\omega t\right)\cos\left(\omega_{c}t+\alpha\right)\hat{J}_{x}\right]e^{-i\omega_{c}t\hat{J}_{z}}-\omega_{c}\hat{J}_{z}$
$=\chi\hat{J}_{z}^{2}+\delta\hat{J}_{z}+2\Omega_{0}\cos\left(\omega t\right)\cos\left(\omega_{c}t+\alpha\right)e^{i\omega_{c}t\hat{J}_{z}}\hat{J}_{x}e^{-i\omega_{c}t\hat{J}_{z}}$
$=\chi\hat{J}_{z}^{2}+\delta\hat{J}_{z}+2\Omega_{0}\cos\left(\omega t\right)\cos\left(\omega_{c}t+\alpha\right)\left[\cos(\omega_{c}t)\hat{J}_{x}-\sin(\omega_{c}t)\hat{J}_{y}\right]$
$=\chi\hat{J}_{z}^{2}+\delta\hat{J}_{z}+\Omega_{0}\cos\left(\omega t\right)\left\{\left[\cos(2\omega_{c}t+\alpha)+\cos\alpha\right]\hat{J}_{x}-\left[\sin(2\omega_{c}t+\alpha)-\sin\alpha\right]\hat{J}_{y}\right\}$
$\approx\chi\hat{J}_{z}^{2}+\delta\hat{J}_{z}+\Omega_{0}\cos\left(\omega t\right)\left(\hat{J}_{x}\cos\alpha+\hat{J}_{y}\sin\alpha\right)$
$=\chi\hat{J}_{z}^{2}+\delta\hat{J}_{z}+\Omega_{0}\cos\left(\omega t\right)\hat{J}_{\alpha},$
where the detuning $\delta=\omega_{0}-\omega_{c}$, and it shows that the Hamiltonian (9) we designed is feasible under near-resonance conditions.
Figure 7: Effects of driving frequency on Floquet-engineered TAT-and-turn for (a) the maximum QFI and (b) the attainable measurement precision via interaction-based readout. Here, $N=100$ and $\phi=1/1000$ are chosen for simulation.
In addition, we also verified that this Floquet-engineered Hamiltonian is equivalent to the TAT-and-turn Hamiltonian under certain conditions. From Fig. 7 we find that when the frequency is large enough, the Floquet-engineered Hamiltonian has the same metrological potential (e.g., QFI) and measurement precision as the ideal TAT-and-turn Hamiltonian. Moreover, under not so high-frequency driving, the measurement precision of interaction-based readout is close to the prediction, which is conducive to experimental implementation.
At last, in real experiments or application scenarios, both the sensing and measurement operators will select $\hat{J}_{z}$, the former for free evolution (phase accumulation) and the latter for half-population readout. We can equivalently perform the desired $\hat{J}_{\vec{n}}$ and $\hat{J}_{\vec{m}}$ with the pulses in the $x-y$ plane. To achieve an equivalent generator $\hat{J}_{\vec{n}}$, assuming the pulses are along $x$-axis and $y$-axis respectively, four pulses are applied before and after the free evolution. Thus we have

### v2

@@ Acknowledgements.
We can further estimate the globally optimal parameter $\delta_{\rm opt}$ and the explicit expression for the time evolution to reach maximum quantum Fisher information (QFI) using the semi-classical description. We find that the TAT-and-turn Hamiltonian can generate GHZ-like states, whose variance is stretched along the $\alpha$-axis or $\beta$-axis on the $x-y$ plane. For calculation clarity, we start from the initial position $(0,0,1)$. When $\delta_{\mathrm{eff}}>0$ (with the default assumption that $\chi>0$), the QFI is given by
$F_{Q}=4\left(\Delta\hat{J}_{\beta}\right)^{2}=N^{2}B^{2}=\frac{N^{2}}{2}(1-Z)(1+Z+\frac{2\delta_{\mathrm{eff}}}{N\chi_{\mathrm{eff}}}),$ (34)
where we take $\hat{J}_{\beta}$ as the generator. When $Z=-\frac{\delta_{\mathrm{eff}}}{N\chi_{\mathrm{eff}}}$, the above Eq. (34) reaches its maximum as a function of $Z$, and it simplifies to
$F_{Q}=\frac{N^{2}}{2}(1+\frac{\delta_{\mathrm{eff}}}{N\chi_{\mathrm{eff}}})^{2}\leq(F_{Q})_{\mathrm{max}}\leq N^{2}.$ (35)
The maximum QFI $(F_{Q})_{\mathrm{max}}$ occurs when
@@
For $\delta_{\mathrm{eff}}<0$, the QFI is expressed as $F_{Q}=4\left(\Delta\hat{J}_{\alpha}\right)^{2}=N^{2}A^{2}\leq N^{2}$ with the generator being $\hat{J}_{\alpha}$. In this case, the optimal detuning is ${\delta_{\rm opt}}/({N\chi_{\mathrm{eff}}})=1-\sqrt{2}\approx-0.414$.
We consider $\delta_{\mathrm{eff}}>0$ (with the analogous calculation for $\delta_{\mathrm{eff}}<0$ provided below), within the time range $0\leq t\leq t_{\rm opt}$ of reaching the maximum QFI. We find that $AB=\frac{1}{2}(1-Z)\sqrt{(1+Z)^{2}-(\frac{2\delta_{\mathrm{eff}}}{N\chi_{\mathrm{eff}}})^{2}}>0$, leading to a differential equation that depends solely on the variable $Z$, which is given by
@@
This result follows from the properties of the uncertainty patch. The lower bound, $\sqrt{1-\frac{1}{2N}}$, of the integral originates from the separatrix of the uncertainty patch, which sets the initial stretching position, while the upper bound, $\frac{2\delta_{\mathrm{eff}}}{N\chi_{\mathrm{eff}}}-1$, corresponds to the critical point of variance change on the separatrix. The Taylor expansion used above is
@@
$f[Z(t)]-f[Z(0)]+\sqrt{{(N\chi_{\mathrm{eff}}})^{2}-\delta_{\mathrm{eff}}^{2}}t=0.$ (42)
@@
Due to the special symmetry of the phase space of the TAT-and-turn Hamiltonian, the same result can be obtained at another saddle point $(0,0,-1)$ for the similar calculation.
III III. Heisenberg scaling and comparison with spin cat states
@@
In the main text, we have shown that the GHZ-like states generated by Floquet-engineered TAT-and-turn dynamics exhibit metrological properties that follow Heisenberg scaling, similar to the spin cat states. In this section, we further analyze the QFI scaling of the GHZ-like states, assuming the generated GHZ-like states can be approximated as spin cat states. Compared to the semi-classical approach, which estimates the critical point $Z(t_{\rm opt})\approx\frac{2\delta_{\mathrm{eff}}}{N\chi_{\mathrm{eff}}}-1$ through Eq. (34), the results obtained here are closer to those from a full quantum treatment.
@@
Applying the time in Eq. (41) and the critical detuning in Eq. (36), the GHZ-like state $\lvert\psi\rangle_{\rm GHZ-like}$ whose two main parts are symmetrically distributed near the pole of $y$-axis can be generated by Floquet-engineered TAT-and-turn dynamics starting from the initial state $\lvert\uparrow\rangle^{\otimes N}$ with $\alpha=0$. To align its orientation with a spin-cat state consisting of two SCSs with the same azimuthal angle $\varphi$ and polar angle symmetric about $\vartheta=\pi/2$, the GHZ-like state $\lvert\psi\rangle_{\rm GHZ-like}$ needs to be rotated to the $z$-axis via $e^{-i\hat{J}_{x}\pi/2}$. The expectation of $\hat{J}_{z}$ for the GHZ-like state $\langle\hat{J}_{z}\rangle_{\rm GHZ-like}^{{}^{\prime}}=0$ with $\lvert\psi\rangle_{\rm GHZ-like}^{{}^{\prime}}=e^{-i\hat{J}_{x}\pi/2}\lvert\psi\rangle_{\rm GHZ-like}$, and the probability distribution is mainly concentrated around the two ends of the $z$-axis. Thus the symmetrical GHZ-like entangle state can be approximately regarded as a macroscopic superposition spin coherent state like a spin cat state [97], which is given by
@@
$\lvert\psi\rangle_{\rm GHZ-like}^{{}^{\prime}}\sim|\Psi(\vartheta)\rangle_{\mathrm{CAT}}\approx\frac{1}{\sqrt{2}}(|\vartheta,\varphi\rangle_{\rm SCS}+|\pi-\vartheta,\varphi\rangle_{\rm SCS}),$ (43)
@@
where $\vartheta=\arccos\overline{Z}$ and the normalized average spin length $\overline{Z}=\frac{2}{N}\sum_{m=-N/2}^{N/2}|m|P_{m}=\frac{2}{N}\sum_{m=-N/2}^{N/2}|m||\langle m|\psi\rangle^{{}^{\prime}}_{\rm{GHZ-like}}|^{2}$ with the eigenstate $|m\rangle$ of $\hat{J}_{z}$. This expression approximates the GHZ-like state as a spin cat state with the same normalized average spin length is $\overline{Z}$. The probability distributions of both the GHZ-like state and its corresponding spin cat state are shown in Fig. 4(a). In this case, the QFI of this GHZ-like state is approximately written as
$F_{Q}^{\rm GHZ-like}\sim F_{Q}^{\rm CAT}\approx N^{2}\cos^{2}\vartheta=N^{2}{\overline{Z}}^{2}.$ (44)
@@
Further, we use Eq. (44) to calculate the QFI of GHZ-like states generated by TAT-and-turn dynamics for different particle number $N$. As shown in Fig. 4(b), the results are in agreement with the quantum results, illustrating that the best achievable QFI of Floquet-engineered TAT-and-turn exhibits the Heisenberg-limited scaling $F_{Q}^{\rm max}\propto N^{2}$, which is consistent with the QCRB of a spin cat state.
Figure 4: (a) The probability distributions of GHZ-like state generated by Floquet-engineered TAT-and-turn dynamics and its corresponding spin cat state with $\theta=\arccos{\overline{Z}}$ for particle number $N=100$. (b) The QFI of GHZ-like states generated by Floquet-engineered TAT-and-turn dynamics and their corresponding spin cat states with $\theta=\arccos{\overline{Z}}$ versus different particle number $N$.
IV IV. Interaction-based readout
@@
The interaction-based readout (IBR) offers a potent protocol for attaining high-precision Heisenberg-limited measurements. When implementing IBR for the final state following the given dynamics evolution, the achievable measurement precision depends on both the directions for encoding the signal into the probe state and for detecting the final state. The signal encoding operator $\hat{J}_{\vec{n}}$ determines the QFI of a given probe state, thus setting a bound (i.e., QCRB) on the ultimate measurement precision. The detection using observable operator $\hat{J}_{\vec{m}}$ determines how closely we can approach the QCRB from a given measurement direction. Optimizing both directions $\vec{n}$ and $\vec{m}$ is crucial for improving measurement precision. In the following, we introduce the method to optimize the measurement direction with a given sensing direction [98, 99].
@@
In the general process of IBR, for an arbitrary observable operator $\hat{J}_{\vec{m}}$, we transform the Schrödinger picture into the Heisenberg picture, i.e.
@@
where $\hat{R}_{\vec{n}}(\phi)=e^{-i\hat{J}_{\vec{n}}\phi}$ for encoding signal, $\hat{U}_{1}=e^{-i\hat{H}_{1}t_{1}}$ for generating entangled states, and $\hat{U}_{2}=e^{-i\hat{H}_{2}t_{2}}$ for signal amplification through disentanglement, typically time-reversal IBR with $\hat{H}_{2}=-\hat{H}_{1}$ and $t_{2}=t_{1}$. When half-population difference readout is performed directly instead of IBR, the Hamiltonian $\hat{H}_{2}=0$. Therefore, the measurement response can be calculated and expressed in matrix form as
@@
where $\mathcal{Q}$ represents the $3\times 3$ covariance matrix with elements $\mathcal{Q}_{\mu\nu}=\langle\{\hat{\tilde{J}}_{\mu},\hat{\tilde{J}}_{\nu}\}\rangle/2-\langle\hat{\tilde{J}}_{\mu}\rangle\langle\hat{\tilde{J}}_{\nu}\rangle$. Using the error propagation formula, the achievable measurement sensitivity can be expressed as
@@
To find the optimal measurement direction, one can apply the Cauchy-Schwarz inequality $(\vec{u}^{T}\vec{v})^{2}\leq\vec{u}^{T}\vec{u}\vec{v}^{T}\vec{v}$, which implies that the measurement sensitivity satisfies
@@
in which $\mathcal{K}=\mathcal{M}\mathcal{Q}^{-1}\mathcal{M}^{T}$ depends on the selection of $\vec{n}$. The equality is achieved at $\vec{m}_{\rm opt}=\mathcal{N}_{c}\mathcal{Q}^{-1}\mathcal{M}^{T}\vec{n}$ with $\mathcal{N}_{c}$ the normalization constant. This applies to a given sensing operator $\hat{J}_{\vec{n}}$ in the general case, including both $\phi=0$ and $\phi\neq 0$.
Thus, by applying the optimal detection direction, i.e. $\vec{m}_{\rm opt}=\mathcal{N}_{c}\mathcal{Q}^{-1}\mathcal{M}^{T}\vec{n}$, the measurement precision is only determined by the encoding signal direction $\vec{n}$. In this work, we choose the sensing direction $\vec{n}_{\rm max}$ that corresponds to the maximum QFI value of the entangled state to encode the signal. By combining Eq. (48) and Eq. (49), we obtain
@@
Note that the phase $\phi$ cannot be equal to 00, as this would cause the matrix $\mathcal{K}$ will lose the direction $\vec{n}_{\rm max}$ and become restricted to the plane perpendicular to the polarization direction.
In real experiments or application scenarios, both the sensing and measurement operators will select $\hat{J}_{z}$, the former for free evolution (phase accumulation) and the latter for half-population readout. We can equivalently perform the desired $\hat{J}_{\vec{n}}$ and $\hat{J}_{\vec{m}}$ with the pulses in the $x-y$ plane. To achieve an equivalent generator $\hat{J}_{\vec{n}}$, assuming the pulses are along $x$-axis and $y$-axis respectively, four pulses are applied before and after the free evolution. Thus we have